A vector R = B + C . If the vector C = 3i + 4 j and vector R has a magnitude equal to that of vector C and it is in positive y direction and what is the magnitude and direction of vector B ?
1
Expert's answer
2011-05-06T05:44:35-0400
Notice that the length of C is √(32+42)=5.
By assumption the length of R is also 5. Since R is in positive y direction, it follows that R=5j, whence B = R-C = (0-3)i + (5-4)j = -3i+j